Hello, I would like to know if someone can help me figure out how to do this type of problem. I'm not so much concerned with getting answers, more I need to understand how to do it myself. Thanks :o)
Solve the system of equations using the substitution method.
3x + y = 2
2x – y = 3

Solve the system of equations using the addition (elimination) method.
3x – 11y = 9
-9x + 33y = 18

Answer by Ashley1516(4)

Hey,
I can help u with the first problem but I'm not sure about the second one.
OK...First, you need to either get the y or the x by itself, with this problem it's easier to get the y in equation 1. Since 3x is positive, subtract it from both sides of the eq. (from itself & 2). Like so:
3x+y=2(eq.1)
-3x -3x
---------
Switch it around:
y=-3x+2 Ah Ha! Now u know what y equals, so u can SUBSTITUTE it into eq.2!
2x-y=3(eq.2)
2x-(-3x+2)=3 Solve this:
2x+3x-2=3 Add the two likes together(x's):
5x-2=3 Add 2 to both sides of the eq.:
+2+2
--------
5x=5 Divide by 5:
*x=1* Ok...U have the x, now just find the y by useing SUBSTITUTION again!
Plug the x into eq.1:
3(1)+y=2 Now solve:
3+y=2 Subtract 3 from both sides:
-3 -3
------
*y=-1* You've got y!
NOW, YOUR ANSWER IS(x,y):
*(1,-1)*
How to CHECK your answer,
Plug in the x & y into both equations:
3x+y=2
3(1)=-1=2
3=-1=2
2=2 CORRECT!
2x-y=3
2(1)-(-1)=2
2-(-1)=3
3=3 CORRECT!
I REALLY hope this helps!
~aShLeY~